A new vaccination is being used in a laboratory experiment to investigate whether it is effective. There are 203 subjects in the study. Is there sufficient evidence to determine if vaccination and disease status are related?

Vaccination Status Diseased Not Diseased Total
Vaccinated 42 37 79
Not Vaccinated 58 66 124
Total 100 103 203

Copy Data

Step 7 of 8:

Make the decision to reject or fail to reject the null hypothesis at the 0.05 level of significance.
Answer #1

Solution:

Test statistic formula is given as below:

Chi square = ∑[(O – E)^2/E]

Where, O is observed frequencies and E is expected frequencies.

E = row total * column total / Grand total

Calculation tables for test statistic are given as below:

Observed Frequencies

Column variable

Row variable

Diseased

Not Diseased

Total

Vaccinated

42

37

79

Not Vaccinated

58

66

124

Total

100

103

203

Expected Frequencies

Column variable

Row variable

Diseased

Not Diseased

Total

Vaccinated

38.91626

40.08374

79

Not Vaccinated

61.08374

62.91626

124

Total

100

103

203

Calculations

(O - E)

3.083744

-3.08374

-3.08374

3.083744

(O - E)^2/E

0.244357

0.23724

0.155679

0.151145

Test Statistic = Chi square = ∑[(O – E)^2/E] = 0.788422

χ2 test statistic = 0.788

Step 7 of 8:

Make the decision to reject or fail to reject the null hypothesis at the 0.05 level of significance.

α = 0.05

We are given

Number of rows = r = 2

Number of columns = c = 2

Degrees of freedom = df = (r – 1)*(c – 1) = 1*1 = 1

Critical value = 3.841459

(by using Chi square table or excel)

P-value = 0.374578

(by using Chi square table or excel)

P-value > α = 0.05

So, we fail to reject the null hypothesis

Answer:

Fail to reject the null hypothesis